A manufacturer of ball bearings claims that this product has a mean weight of 5.02 g and a standard deviation of 0.30 g. What is the probability that a random sample of 100 ball bearings will have combined weight:


a. between 496 g and 500 g?


b. more than 510 g?


1
Expert's answer
2021-03-30T16:00:41-0400

X ~ N(5.02, 0.30^2)

n = 100

a.

"P(4.96< \bar{X}<5.00) = P(\frac{4.96-5.02}{\frac{0.3}{\sqrt{100}}} <Z< \frac{5.00-5.02}{\frac{0.3}{\sqrt{100}}}) \\

= P(\frac{4.96-5.02}{\frac{0.3}{\sqrt{100}}} <Z< \frac{5.00-5.02}{\frac{0.3}{\sqrt{100}}}) \\

= P(-2.00 < Z < -0.666) \\

= P(Z< -0.666) -P(Z< -2) \\

= 0.2527 -0.0227 \\

= 0.2300 \\

= 23.00 \; \%"

b.

"P(\bar{X}>5.10) = 1 -P(\bar{X}<5.1) \\

= 1 -P(Z< \frac{x-μ}{\frac{σ}{\sqrt{n}}}) \\

= 1 - P(Z< \frac{5.10-5.02}{\frac{0.3}{\sqrt{100}}}) \\

= 1 -P(Z<2.666) \\

= 1 -0.9961 \\

= 0.0039 \\

= 0.39 \; \%"